OpenStudy (anonymous):

If the Wronskian W of f and g is 3e^(4t) and if f(t)=e^(2t), find g(t).

OpenStudy (anonymous):

Since the Wronskian is \[W(f,g)=fg'-gf'=3e^{4t}\] and \[f(t)=e^{2t}\text{ AND } f'(t)=2e^{2t}\] then \[g'e^{2t}-2ge^{2t}=3e^{4t}\] Divide by $e^{2t}$ \[g'-2g=3e^{2t}\] This is a nonhomogenous first order differential equation. We write this in the form \[g'+pg=h\] There is a really easy way to solve this. We introduce a dummy factor called the integrating factor. Let \[u'(t)=u(t)p\] \[ug'+gu'=(gu)'=uh\] and \[g=\frac{\int uh dt}{u}\] The integrating factor is \[u(t)=e^{\int- 2dt+c} \] \[u(t)=e^{-2t+c} \] Then, \[g(t)=\frac{\int u(t)(3e^{2t})dt+k}{u(t)}=\frac{\int e^{-2t+c}3e^{2t}dt+k}{e^{-2t+c}}=\frac{3\int e^{c}dt+k}{e^{-2t+c}}=\frac{3te^c+k}{e^{-2t+c}}\] So \[g(t)=3te^{2t}+ke^{2t}\]

OpenStudy (anonymous):

You are very detailed, and really really helped me understand what I was supposed to be doing. I thank you for that.